# FreeInsert: Disentangled Text-Guided Object Insertion in 3D Gaussian Scene without Spatial Priors

### v3

@@ 3.5. Object Appearance Refinement
Once the object’s degrees of freedom are determined, a refinement module is introduced to enhance the visual quality of the inserted object $\mathcal{G}_{O}$. Specifically, we refine $\mathcal{G}_{O}$ using the high-quality appearance from the inserted-object image $\mathcal{I}_{O}$ via LoRA (Hu et al., 2022).
Viewpoint Frequency Balancing. To avoid the overfitting caused by using a single-view optimization, which often leads to 3D inconsistencies and missing object parts (e.g., a side view causing missing legs) as shown in Figure 10 (b). We perform multi-view sampling of the inserted object. Specifically, given a set of views $\{I_{i}\}_{i=1}^{N}$, rendered from the inserted object $\mathcal{G}_{O}$. We estimate the pose $P^{*}$ of the object image $\mathcal{I}_{\textit{O}}$ by selecting the most similar view based on DINO feature similarity (Oquab et al., 2024). To ensure both appearance fidelity and geometric consistency, We construct the training set $\mathcal{D}_{\text{ref}}$ by combining the rendered multi-view images with repeated samples of the inserted-object image $\mathcal{I}_{\textit{O}}$ and its estimated pose $P^{*}$, as follows:
@@
$z_{i}$ is the noisy latent of image $I_{i}$, $t$ is the diffusion timestep, and $\epsilon$ is the target noise. The denoising network $\epsilon_{\phi_{2}}$, augmented with LoRA (Hu et al., 2022), is conditioned on $I_{i}$, its pose $P_{i}$, and the object-specific prompt $y^{*}$, e.g., “A <token> dog”, which is formatting from $\mathcal{T}_{\textit{O}}$.
Appearance-Focused Refinement. We employ fine-tuning diffusion to update the object Gaussian $\mathcal{G}_{O}$, guided by $\mathcal{L}_{sds}$ (Poole et al., 2022). Under our setting, we sample from a lower range of timesteps during optimization to reduce the impact on the object’s geometry (e.g., shape and scale), thereby encouraging the model to focus more on refining appearance details. The corresponding objective is defined as follows:
@@ 4. Experiments
Dataset To comprehensively evaluate our method, we follow prior works (Chen et al., 2024a; Zhuang et al., 2024; Haque et al., 2023) and select representative scenes of varying complexity, including simple backgrounds, human faces, and complex outdoor environments. In these scenes, we insert commonly associated objects (e.g., glasses, giraffes) and evaluate diverse categories such as bowties and moustaches to assess generalization. For GaussianEditor (Chen et al., 2024a), we manually annotate masks, while for TIP-Editor (Zhuang et al., 2024), we use the author-provided bounding boxes and object images for comparison.
Baselines We compare our method with state-of-the-art 3D scene editing approaches that support object insertion and replacement, under two types of guidance: text prompt and text-image prompt.The text-guided baselines include three methods: Instruct-GS2GS (Vachha and Haque, [n. d.]), which extends Instruct-NeRF2NeRF (IN2N) (Haque et al., 2023) by replacing the NeRF in IN2N with a 3DGS model; GaussCtrl (Wu et al., 2024), and GaussianEditor (Chen et al., 2024a). For text-image prompt methods, we compare with TIP-Editor (Zhuang et al., 2024), which uses an example image to specify object appearance. As TIP-Editor provides only limited insertion scripts (e.g., “A doll wearing sunglasses”, “A man with beard”). For fairness, we use official code and pre-trained weights.
@@
Evaluation Criteria. We use CLIP Text-Image directional similarity following (Chen et al., 2024a; Zhuang et al., 2024; Wu et al., 2024; Zhuang et al., 2023) to assess the alignment between the text and the editing results. For appearance-specified cases, we further employ DINO similarity (Oquab et al., 2024) following (Zhuang et al., 2024) to assess appearance preservation. We also conducted a user study with 50 participants, who rated the 3D editing results (presented with prompts in shuffled order) on four criteria: Semantic Alignment, Object Integrity, Geometric Consistency, and Detail Preservation, using a 1–10 scale.
@@ 4.2.1. Qualitative comparisons
Text-Image Prompt Comparisons. Besides the text-prompt methods, we further evaluate object insertion and replacement with a given image prompt of the specified object in Figure 4, comparing against TIP-Editor (Zhuang et al., 2024). Although TIP-Editor supports flexible insertion via 3D bounding boxes, it suffers from inconsistent multi-view appearances due to its reliance on 2D editing techniques. Most critically, achieving such results remains dependent on finely user-provided 3D bounding boxes, which significantly hinders scalability and practicality. In contrast, our method delivers the most complete geometry and better appearance fidelity to the image prompt without relying on any annotation.
@@
InstructN2N(GS) (Vachha and Haque, [n. d.]) 26.76% - 18.3
GaussianEditor (Chen et al., 2024a) 27.36% - 24.7
GaussCtrl (Wu et al., 2024) 25.39% - 26.3
TIP-Editor (Zhuang et al., 2024) 30.01% 83.30% 32.3
@@ 4.3. Ablation Studies
Comparison of DoF learning directly by Different Multi-Modal Large Language Models vs. FreeInsert. To evaluate our DoF learning method, we compare it with state-of-the-art multimodal LLMs, including GPT-4V (Achiam et al., 2023), Molmo-7B (Deitke et al., 2024), and GPT-o1 (Jaech et al., 2024), which directly predict object DoFs. Translation and scale are derived from multi-view prompts (e.g., “Point the four coordinates of a bounding box to add [object] to/on [target]”) and lifted to 3D, while rotation follows our initialization. As shown in Figure 9 (“Add a pair of glasses to the doll”), our predictions are more plausible. Quantitatively, our method achieves the highest alignment with human preferences in projected mIoU across all cases Table 3.
GPT-4V(Achiam et al., 2023)
Molmo-7B(Deitke et al., 2024)
GPT-o1(Jaech et al., 2024)

### v4

@@ 3.5. Object Appearance Refinement
Once the object’s degrees of freedom are determined, a refinement module is introduced to enhance the visual quality of the inserted object $\mathcal{G}_{O}$. Specifically, we refine $\mathcal{G}_{O}$ using the high-quality appearance from the inserted-object image $\mathcal{I}_{O}$ via LoRA (hu2022lora).
Viewpoint Frequency Balancing. To avoid the overfitting caused by using a single-view optimization, which often leads to 3D inconsistencies and missing object parts (e.g., a side view causing missing legs) as shown in Figure 10 (b). We perform multi-view sampling of the inserted object. Specifically, given a set of views $\{I_{i}\}_{i=1}^{N}$, rendered from the inserted object $\mathcal{G}_{O}$. We estimate the pose $P^{*}$ of the object image $\mathcal{I}_{\textit{O}}$ by selecting the most similar view based on DINO feature similarity (oquab2024dinov). To ensure both appearance fidelity and geometric consistency, We construct the training set $\mathcal{D}_{\text{ref}}$ by combining the rendered multi-view images with repeated samples of the inserted-object image $\mathcal{I}_{\textit{O}}$ and its estimated pose $P^{*}$, as follows:
@@
$z_{i}$ is the noisy latent of image $I_{i}$, $t$ is the diffusion timestep, and $\epsilon$ is the target noise. The denoising network $\epsilon_{\phi_{2}}$, augmented with LoRA (hu2022lora), is conditioned on $I_{i}$, its pose $P_{i}$, and the object-specific prompt $y^{*}$, e.g., “A <token> dog”, which is formatting from $\mathcal{T}_{\textit{O}}$.
Appearance-Focused Refinement. We employ fine-tuning diffusion to update the object Gaussian $\mathcal{G}_{O}$, guided by $\mathcal{L}_{sds}$ (poole2022dreamfusiontextto3dusing2d). Under our setting, we sample from a lower range of timesteps during optimization to reduce the impact on the object’s geometry (e.g., shape and scale), thereby encouraging the model to focus more on refining appearance details. The corresponding objective is defined as follows:
@@ 4. Experiments
Dataset To comprehensively evaluate our method, we follow prior works (chen2024gaussianeditor; zhuang2024tip; haque2023instruct) and select representative scenes of varying complexity, including simple backgrounds, human faces, and complex outdoor environments. In these scenes, we insert commonly associated objects (e.g., glasses, giraffes) and evaluate diverse categories such as bowties and moustaches to assess generalization. For GaussianEditor (chen2024gaussianeditor), we manually annotate masks, while for TIP-Editor (zhuang2024tip), we use the author-provided bounding boxes and object images for comparison.
Baselines We compare our method with state-of-the-art 3D scene editing approaches that support object insertion and replacement, under two types of guidance: text prompt and text-image prompt.The text-guided baselines include three methods: Instruct-GS2GS (vachhainstruct), which extends Instruct-NeRF2NeRF (IN2N) (haque2023instruct) by replacing the NeRF in IN2N with a 3DGS model; GaussCtrl (wu2024gaussctrl), and GaussianEditor (chen2024gaussianeditor). For text-image prompt methods, we compare with TIP-Editor (zhuang2024tip), which uses an example image to specify object appearance. As TIP-Editor provides only limited insertion scripts (e.g., “A doll wearing sunglasses”, “A man with beard”). For fairness, we use official code and pre-trained weights.
@@
Evaluation Criteria. We use CLIP Text-Image directional similarity following (chen2024gaussianeditor; zhuang2024tip; wu2024gaussctrl; zhuang2023dreameditor) to assess the alignment between the text and the editing results. For appearance-specified cases, we further employ DINO similarity (oquab2024dinov) following (zhuang2024tip) to assess appearance preservation. We also conducted a user study with 50 participants, who rated the 3D editing results (presented with prompts in shuffled order) on four criteria: Semantic Alignment, Object Integrity, Geometric Consistency, and Detail Preservation, using a 1–10 scale.
@@ 4.2.1. Qualitative comparisons
Text-Image Prompt Comparisons. Besides the text-prompt methods, we further evaluate object insertion and replacement with a given image prompt of the specified object in Figure 4, comparing against TIP-Editor (zhuang2024tip). Although TIP-Editor supports flexible insertion via 3D bounding boxes, it suffers from inconsistent multi-view appearances due to its reliance on 2D editing techniques. Most critically, achieving such results remains dependent on finely user-provided 3D bounding boxes, which significantly hinders scalability and practicality. In contrast, our method delivers the most complete geometry and better appearance fidelity to the image prompt without relying on any annotation.
@@
InstructN2N(GS) (vachhainstruct) 26.76% - 18.3
GaussianEditor (chen2024gaussianeditor) 27.36% - 24.7
GaussCtrl (wu2024gaussctrl) 25.39% - 26.3
TIP-Editor (zhuang2024tip) 30.01% 83.30% 32.3
@@ 4.3. Ablation Studies
Comparison of DoF learning directly by Different Multi-Modal Large Language Models vs. FreeInsert. To evaluate our DoF learning method, we compare it with state-of-the-art multimodal LLMs, including GPT-4V (achiam2023gpt), Molmo-7B (deitke2024molmo), and GPT-o1 (jaech2024openai), which directly predict object DoFs. Translation and scale are derived from multi-view prompts (e.g., “Point the four coordinates of a bounding box to add [object] to/on [target]”) and lifted to 3D, while rotation follows our initialization. As shown in Figure 9 (“Add a pair of glasses to the doll”), our predictions are more plausible. Quantitatively, our method achieves the highest alignment with human preferences in projected mIoU across all cases Table 3.
GPT-4V(achiam2023gpt)
Molmo-7B(deitke2024molmo)
GPT-o1(jaech2024openai)